# with LLM-Generated Documents Integration

### v1

@@ 1 Introduction
Figure 1: Ranking performance versus source bias comparison on averaged results of 16 datasets benchmarked in Cocktail. A more negative Relative $\Delta$ signifies increased source bias towards LLM-generated content. The Pearson correlation coefficient between these two axes is $-0.798$ ($p$-value < 0.05), indicating a strong negative correlation. For brevity, we omit the ‘%’ symbol of the scores in all the tables and figures.
@@ 3.2 Dataset Statistics and Analysis
Following practices in Dai et al. (2024b), our dataset construction process involved feeding only the original human-written text to the LLM for rewriting, without inputting any query-specific information to avoid introducing additional query-related bias. Furthermore, we employed the OpenAI embedding model111text-embedding-ada-002: https://platform.openai.com/docs/guides/embeddings to obtain semantic embeddings for both text sources, comparing them through cosine similarity. These comparisons, presented in Appendix Figure 7, demonstrate a high similarity, indicating successful semantic retention of the original human-written text in the LLM-generated texts. Additionally, evaluations of several retrieval models on sole human-written or LLM-generated corpora showed consistent performance, as seen in Appendix Table 8. Moreover, we also conduct human evaluations for quality verification for each dataset in Appendix Table 10. These observations reinforce confidence in the quality of our newly constructed datasets, suggesting that the LLM-generated content maintains semantics comparable to human-written texts for IR tasks. The detailed dataset analysis is provided in AppendixA.3.
@@
TREC-COVID 67.0 67.0 62.0 68.0 65.0 64.0 70.0 72.0 75.0 77.0 85.0 70.2 70.5
NFCorpus 45.4 39.2 36.4 35.3 41.5 42.6 43.5 41.6 42.7 50.2 49.4 42.5 42.2
NQ 45.7 62.5 60.0 60.0 65.0 69.6 65.6 67.5 70.4 69.1 68.9 64.0 65.9
HotpotQA 84.3 75.3 62.9 72.1 84.6 88.1 82.1 88.5 89.1 93.9 90.7 82.9 82.7
FiQA-2018 24.4 24.7 24.9 28.9 28.1 34.0 27.2 31.8 36.3 35.5 38.7 30.4 31.0
@@
FEVER 65.6 80.6 75.9 80.4 83.1 86.4 75.8 87.6 88.0 90.2 43.4 77.9 79.1
Climate-FEVER 25.2 26.4 27.4 28.3 32.7 31.4 28.1 32.3 34.0 35.2 35.3 30.6 31.1
SciFact 55.7 38.0 40.3 40.7 53.7 55.0 50.0 52.3 55.3 56.3 65.0 51.1 50.7
@@
Supervised 50.6 69.9 68.4 66.2 66.2 67.2 70.3 69.9 73.3 70.8 69.7 67.5 69.2
Zero-shot 48.4 47.9 45.3 48.6 52.0 53.9 50.5 54.0 56.4 58.9 54.6 51.8 52.2
All 48.8 52.0 49.6 51.9 54.7 56.4 54.2 57.0 59.6 61.1 57.4 54.8 55.4
@@ 4 Benchmarking Evaluation Protocol
Evaluation Framework. To standardize the assessment of IR models within the LLM era, we also develop a Python framework that combines user-friendliness with comprehensive evaluation capabilities. Built on the foundation of the BEIR Thakur et al. (2021), our framework inherits its best features, including the ability to easily replicate experiments from open-sourced repositories and incorporate new models and datasets. A key innovation of our framework is its ability to conduct evaluations using either individual or mixed corpora, accommodating the mixture of human-written and LLM-generated content characteristic of the LLM era. These features make our framework an invaluable tool for advancing IR research and application in both academia and industry.
@@ 5.2 Benchmarked Results
NDCG@1 48.8 61.1 57.4 59.6 $\uparrow$10.8 59.8 $\downarrow$1.3 58.5 $\uparrow$1.1
@@
Bias in the first retrieval stage tends to propagate and even amplify during the re-ranking stage. This trend is particularly notable in datasets like NFCorpus and Touché-2020, where the source bias observed in the first retrieval stage persists and tends to intensify during the second-stage re-ranking. Moreover, as detailed in Table 4, enhancing the efficiency of the first-stage retrieval by replacing BM25 with DRAGON does not necessarily improve performance in the subsequent re-ranking phase. However, the source bias inherent in the first-stage retriever significantly impacts and may even magnify in subsequent re-ranking. This observation underscores the critical need for developing holistic approaches to mitigate bias throughout the retrieval pipeline, ensuring fairness and accuracy in the whole IR system.
@@ Acknowledgements
Jun Xu is the corresponding author. This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No. 62377044, No.62376275), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China. This work was partially done at Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education.
@@ A.1.4 BEIR
If not specified, we utilize the preprocessed version of the following datasets as included in the BEIR benchmark Thakur et al. (2021) to serve as the human-written seed datasets in our Cocktail benchmark, ensuring consistency and comparability in our evaluations.
@@ A.2 Dataset Processing and Quality Control
In refining each corpus, we only kept documents with text lengths between 10 and 2,000 words. This criterion was set because texts shorter than 10 words often consisted of empty texts or symbols, compromising data quality—this was particularly evident in the TREC-COVID dataset, which contained over 40,000 documents with empty text. Conversely, texts longer than 2,000 words were filtered out to accommodate the context length limitation of 4,096 tokens for rewriting with Llama2, impacting only 0.05% of the data. These filters did not significantly alter the overall data distribution, with a total of less than 1% data removed, and our evaluation results align closely with those reported in the BEIR benchmark Thakur et al. (2021) for the original datasets, as shown in Table 2.
When LLM is used for rewriting, LLM will inevitably refuse to rewrite due to the safety constraints, especially for datasets like Touché-2020, which focus on retrieving contentious viewpoints. For such LLM-generated data, we keep the same content as the human-written counterparts. This approach allows for the easy removal of refused rewrites if necessary, ensuring both the integrity and the quality of the dataset processing and control measures implemented. The statistics of the final 16 benchmarked datasets in Cocktails are summarized in Table 1.
@@ A.3 Dataset Statistics and Analysis
Quality Verification with Human Evaluation. To further validate the assigned relevancy label, we also conduct a human evaluation study. Due to cost constraints of human evaluation, we sample 20 <query, human doc, LLM doc> triples from each of the 16 datasets included in Cocktail. These triples were annotated by graduate students and our colleagues, who were asked to assess which document is more semantically relevant to the given query, with options being “human doc”, “LLM doc”, or “equal”. At least three different annotators evaluated each triple, with the majority vote deciding the final label. We summarized the human evaluation results in Table 10, with numbers in parentheses indicating the percentage of agreement among all three evaluators for each option. The results show a comparable level of semantic relevance between human-written and LLM-generated texts for the given queries, ensuring the fairness of our analysis of source bias.
@@ B.2 Checkpoints and Implementation Details
For the fine-tuning of BERT and RoBERTa models, we utilized the official training script [^5] https://github.com/beir-cellar/beir/blob/main/examples/retrieval/training/train_msmarco_v3.py provided in the BEIR benchmark on the MSMARCO training dataset. Unless otherwise noted, mean pooling was employed as the pooling strategy, alongside cosine similarity as the scoring function. Training parameters for each model included a duration of $10$ epochs, utilizing a batch size of $75$, a learning rate set at $-52e\text{-}$ 5, and the AdamW optimizer for efficient optimization. Our software framework utilizes PyTorch version 2.0.0 and the HuggingFace Transformers library version 4.31.0. All experiments were performed using approximately 100 hours on four NVIDIA RTX A6000 (48G) GPUs.
@@ Appendix C More Experimental Results
TREC-COVID 62.4 66.4 60.0 71.0 64.0 63.2 70.9 74.3 68.3 71.9 82.1 68.6 69.2
NFCorpus 44.2 36.3 32.3 33.0 39.1 42.1 41.1 39.3 41.1 48.2 48.0 40.4 40.1
NQ 46.2 62.5 60.0 60.3 66.3 70.6 66.0 67.9 71.2 68.6 69.1 64.4 66.3
HotpotQA 72.3 62.3 51.7 59.2 72.3 76.3 69.4 76.5 77.0 83.5 81.4 71.1 71.0
FiQA-2018 22.2 21.9 22.2 26.2 26.1 29.9 25.3 28.6 33.5 31.6 34.9 27.5 28.0
Touché-2020 51.4 47.0 42.1 46.8 44.7 46.3 34.3 47.1 53.4 52.8 52.5 47.1 46.7
CQADupStack 26.4 24.8 25.0 29.5 23.3 32.6 31.7 29.5 34.8 32.6 34.4 29.5 29.8
DBPedia 34.2 52.6 47.3 45.8 54.5 59.3 54.8 55.1 57.7 57.4 33.2 50.2 51.8
SCIDOCS 15.4 11.7 10.5 12.7 15.0 15.5 13.9 14.8 16.1 17.4 18.3 14.7 14.6
FEVER 67.5 80.5 75.8 80.7 83.8 86.5 75.5 87.7 87.9 89.1 44.6 78.1 79.2
Climate-FEVER 23.4 25.1 25.2 25.9 30.3 29.6 25.4 29.9 31.0 32.7 33.0 28.3 28.8
SciFact 58.3 37.9 40.9 40.8 54.0 56.4 49.9 54.5 56.7 57.1 65.4 52.0 51.4
@@
Supervised 47.4 68.2 68.2 65.4 68.3 67.4 69.8 70.3 72.2 71.6 69.7 67.1 69.1
Zero-shot 45.7 46.3 43.1 46.6 50.2 52.6 48.3 52.4 54.3 56.1 52.5 49.8 50.2
All 46.0 50.4 47.8 50.2 53.6 55.4 52.3 55.7 57.6 59.0 55.8 53.1 53.8
@@
TREC-COVID 61.3 63.2 57.7 68.1 64.4 62.0 70.7 74.4 68.4 70.9 80.1 67.4 68.0
NFCorpus 41.5 34.3 30.0 30.1 37.5 39.7 38.2 37.3 39.7 45.4 45.6 38.1 37.8
NQ 49.3 64.8 62.3 62.8 68.9 73.5 68.7 70.4 73.7 70.2 70.9 66.9 68.6
HotpotQA 67.8 56.6 46.8 53.7 67.4 71.4 64.0 71.6 72.0 79.0 77.8 66.2 66.0
FiQA-2018 21.6 21.1 22.1 25.5 25.7 29.2 24.4 27.8 32.3 30.5 33.7 26.7 27.2
Touché-2020 50.4 45.6 40.9 48.7 42.9 43.7 32.6 44.4 50.1 50.2 52.0 45.6 45.1
CQADupStack 27.3 25.4 25.7 30.1 24.5 33.5 32.6 30.2 35.7 33.1 34.8 30.3 30.6
DBPedia 33.3 49.2 44.4 43.5 51.8 55.5 52.6 52.6 54.0 55.1 33.3 47.8 49.2
SCIDOCS 14.1 10.7 9.5 11.4 13.3 14.2 12.4 13.6 14.8 15.8 16.5 13.3 13.2
FEVER 71.4 82.5 78.1 82.4 85.8 88.1 78.5 89.2 89.5 89.9 48.1 80.3 81.2
Climate-FEVER 22.1 23.8 23.9 24.0 28.6 28.0 23.7 27.9 29.0 31.3 31.5 26.7 27.2
SciFact 62.0 41.0 43.0 42.9 56.4 59.6 53.0 57.1 59.9 61.2 68.2 54.9 54.2
@@
Supervised 47.2 69.4 68.5 66.7 69.9 68.5 70.7 70.8 73.7 71.3 69.4 67.8 69.9
Zero-shot 45.4 45.3 42.2 45.8 49.5 51.7 47.8 51.6 53.4 55.1 52.1 49.1 49.4
All 45.8 49.9 47.1 49.7 53.3 54.9 52.1 55.2 57.2 58.2 55.3 52.6 53.3

### v2

@@ 1 Introduction
Figure 1: Ranking performance versus source bias comparison on averaged results of 16 datasets benchmarked in Cocktail. A more negative Relative $\Delta$ signifies increased source bias towards LLM-generated content. The Pearson correlation coefficient between these two axes is $-0.772$ ($p$-value < 0.05), indicating a strong negative correlation. For brevity, we omit the ‘%’ symbol of the scores in all the tables and figures.
@@ 3.2 Dataset Statistics and Analysis
Following practices in Dai et al. (2024b), our dataset construction process involved feeding only the original human-written text to the LLM for rewriting, without inputting any query-specific information to avoid introducing additional query-related bias. Furthermore, we employed the OpenAI embedding model111text-embedding-ada-002: https://platform.openai.com/docs/guides/embeddings to obtain semantic embeddings for both text sources, comparing them through cosine similarity. For calibrating the results, we also compare the semantic similarity between randomly selected <human-written doc, LLM-generated doc> pairs and matching pairs <human-written doc, LLM-generated doc>. These comparisons, presented in Appendix Figure 7 and Table 9, demonstrate a high similarity between the LLM-generated document and the corresponding original human-written counterpart, indicating successful semantic retention of the original human-written text in the LLM-generated texts. Additionally, evaluations of several retrieval models on sole human-written or LLM-generated corpora showed consistent performance, as seen in Appendix Table 8. Moreover, we also conduct human evaluations for quality verification for each dataset in Appendix Table 10. These observations reinforce confidence in the quality of our newly constructed datasets, suggesting that the LLM-generated content maintains semantics comparable to human-written texts for IR tasks. The detailed dataset analysis is provided in Appendix A.3.
@@
TREC-COVID 67.0 67.0 62.0 68.0 65.0 64.0 70.0 72.0 75.0 77.0 82.0 69.9 70.2
NFCorpus 45.4 39.2 36.4 35.3 41.5 42.6 43.5 41.6 42.7 50.2 44.1 42.0 41.7
NQ 45.7 62.5 60.0 60.0 65.0 69.6 65.6 67.5 70.4 69.1 71.3 64.2 66.1
HotpotQA 84.3 75.3 62.9 72.1 84.6 88.1 82.1 88.5 89.1 93.9 89.8 82.8 82.6
FiQA-2018 24.4 24.7 24.9 28.9 28.1 34.0 27.2 31.8 36.3 35.5 40.9 30.6 31.2
@@
FEVER 65.6 80.6 75.9 80.4 83.1 86.4 75.8 87.6 88.0 90.2 76.3 80.9 82.4
Climate-FEVER 25.2 26.4 27.4 28.3 32.7 31.4 28.1 32.3 34.0 35.2 33.8 30.4 31.0
SciFact 55.7 38.0 40.3 40.7 53.7 55.0 50.0 52.3 55.3 56.3 65.3 51.2 50.7
@@
Supervised 50.6 69.9 68.4 66.2 66.2 67.2 70.3 69.9 73.3 70.8 66.8 67.2 68.9
Zero-shot 48.4 47.9 45.3 48.6 52.0 53.9 50.5 54.0 56.4 58.9 58.0 52.1 52.5
All 48.8 52.0 49.6 51.9 54.7 56.4 54.2 57.0 59.6 61.1 59.7 55.0 55.6
@@ 4 Benchmarking Evaluation Protocol
Evaluation Framework. To standardize the assessment of IR models within the LLM era, we also develop a Python framework that combines user-friendliness with comprehensive evaluation capabilities. Built on the foundation of the BEIR Thakur et al. (2021), our framework inherits its best features, including the ability to easily replicate experiments from open-sourced repositories and incorporate new models and datasets. A key innovation of our framework is its ability to conduct evaluations using either individual or mixed corpora, accommodating the mixture of human-written and LLM-generated content characteristic of the LLM era. These features make our framework an invaluable tool for advancing IR research and application in both academia and industry. A detailed description and all the scripts of our evaluation framework are provided in the link https://github.com/KID-22/Cocktail.
@@ 5.2 Benchmarked Results
NDCG@1 48.8 61.1 59.7 59.6 $\uparrow$10.8 61.0 $\downarrow$0.1 59.7 $\downarrow$0.0
@@
Bias in the first retrieval stage tends to propagate and even amplify during the re-ranking stage. This trend is particularly notable in datasets like NFCorpus and Touché-2020, where the source bias observed in the first retrieval stage persists and tends to intensify during the second-stage re-ranking. Moreover, as detailed in Table 4, enhancing the efficiency of the first-stage retrieval by replacing BM25 with DRAGON does not necessarily improve performance in the subsequent re-ranking phase. However, the source bias inherent in the first-stage retriever significantly impacts and may even magnify in subsequent re-ranking. This observation underscores the critical need for developing holistic approaches to mitigate bias throughout the retrieval pipeline, thereby ensuring fairness and accuracy in the whole IR system.
@@ Acknowledgements
This work was funded by the National Key R&D Program of China (2023YFA1008704), the National Natural Science Foundation of China (No. 62377044, No.62276248), Beijing Key Laboratory of Big Data Management and Analysis Methods, Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class” Initiative, PCC@RUC, funds for building world-class universities (disciplines) of Renmin University of China, and the Youth Innovation Promotion Association CAS under Grants No.2023111. This work was supported by the Fundamental Research Funds for the Central Universities, and the Research Funds of Renmin University of China (RUC24QSDL013).
@@ A.1.4 BEIR
If not specified, we utilize the preprocessed version of the following open-sourced datasets as included in the BEIR benchmark Thakur et al. (2021) to serve as the human-written seed datasets in our Cocktail benchmark, ensuring consistency and comparability in our evaluations.
@@ A.2 Dataset Processing and Quality Control
Documents with text lengths between 10 and 2,000 words are retained for each corpus. This criterion was set because texts shorter than 10 words often consisted of empty texts or symbols, compromising data quality—this was particularly evident in the TREC-COVID dataset, which contained over 40,000 documents with empty text. Conversely, texts over 2,000 words were filtered out to satisfy the context length limitation of 4,096 tokens for rewriting with Llama2, impacting only 0.05% of the data. These filters did not significantly alter the overall data distribution, with a total of less than 1% data removed, and our evaluation results align closely with those reported in the BEIR benchmark Thakur et al. (2021) for the original datasets, as shown in Table 2. For most datasets, we generated LLM Corpus once. However, for MS MARCO, HotpotQA, DBPedia, and NQ-UTD, to construct a more challenging benchmark for evaluating source bias in IR models, we generated multiple versions and used BERT to filter and merge the corpus, aiming to increase the task difficulty.
When LLM is used for rewriting, LLM will inevitably refuse to rewrite due to safety constraints, especially for datasets like Touché-2020, which focus on retrieving contentious viewpoints. For such LLM-generated data, we keep the same content as the human-written counterparts. This approach allows for the easy removal of refused rewrites if necessary, ensuring both the integrity and the quality of the dataset processing and control measures implemented. The statistics of the 16 benchmarked datasets in Cocktails are summarized in Table 1.
@@ A.3 Dataset Statistics and Analysis
Quality Verification with Human Evaluation. To further validate the assigned relevancy label, we also conduct a human evaluation study. Due to cost constraints of human evaluation, we sample 20 <query, human doc, LLM doc> triples from each of the 16 datasets included in Cocktail. These triples were annotated by graduate students and our colleagues, who were asked to assess which document is more semantically relevant to the given query, with options being “human doc”, “LLM doc”, or “equal”. At least three different annotators evaluated each triple, with the majority vote deciding the final label. We summarized the results in Table 10, with numbers in parentheses indicating the percentage of agreement among all three evaluators for each option. The results show a comparable level of semantic relevance between human-written and LLM-generated texts for the given queries, ensuring the fairness of our analysis of source bias.
@@ B.2 Checkpoints and Implementation Details
For the fine-tuning of BERT and RoBERTa models, we utilized the official training script [^5] https://github.com/beir-cellar/beir/blob/main/examples/retrieval/training/train_msmarco_v3.py in BEIR on the MSMARCO training dataset. Unless otherwise noted, mean pooling was employed as the pooling strategy, alongside cosine similarity as the scoring function. Training parameters for each model included a duration of $10$ epochs, utilizing a batch size of $75$, a learning rate set at $-52e\text{-}$ 5, and the AdamW optimizer for efficient optimization. Our software framework utilizes PyTorch version 2.0.0 and the HuggingFace Transformers library version 4.31.0. All experiments were performed using approximately 100 hours on four NVIDIA RTX A6000 (48G) GPUs.
@@ Appendix C More Experimental Results
TREC-COVID 62.4 66.4 60.0 71.0 64.0 63.2 70.9 74.3 68.3 71.9 79.9 68.4 69.0
NFCorpus 44.2 36.3 32.3 33.0 39.1 42.1 41.1 39.3 41.1 48.2 43.1 40.0 39.6
NQ 46.2 62.5 60.0 60.3 66.3 70.6 66.0 67.9 71.2 68.6 72.7 64.7 66.6
HotpotQA 72.3 62.3 51.7 59.2 72.3 76.3 69.4 76.5 77.0 83.5 80.1 71.0 70.8
FiQA-2018 22.2 21.9 22.2 26.2 26.1 29.9 25.3 28.6 33.5 31.6 37.3 27.7 28.3
Touché-2020 51.4 47.0 42.1 46.8 44.7 46.3 34.3 47.1 53.4 52.8 51.7 47.1 46.6
CQADupStack 26.4 24.8 25.0 29.5 23.3 32.6 31.7 29.5 34.8 32.6 35.2 29.6 29.9
DBPedia 34.2 52.6 47.3 45.8 54.5 59.3 54.8 55.1 57.7 57.4 49.6 51.7 53.4
SCIDOCS 15.4 11.7 10.5 12.7 15.0 15.5 13.9 14.8 16.1 17.4 18.4 14.7 14.6
FEVER 67.5 80.5 75.8 80.7 83.8 86.5 75.5 87.7 87.9 89.1 77.0 81.1 82.4
Climate-FEVER 23.4 25.1 25.2 25.9 30.3 29.6 25.4 29.9 31.0 32.7 31.8 28.2 28.7
SciFact 58.3 37.9 40.9 40.8 54.0 56.4 49.9 54.5 56.7 57.1 65.2 52.0 51.3
@@
Supervised 47.4 68.2 68.2 65.4 68.3 67.4 69.8 70.3 72.2 71.6 69.0 67.1 69.0
Zero-shot 45.7 46.3 43.1 46.6 50.2 52.6 48.3 52.4 54.3 56.1 56.1 50.1 50.6
All 46.0 50.4 47.8 50.2 53.6 55.4 52.3 55.7 57.6 59.0 58.5 53.3 54.0
@@
TREC-COVID 61.3 63.2 57.7 68.1 64.4 62.0 70.7 74.4 68.4 70.9 80.0 67.4 68.0
NFCorpus 41.5 34.3 30.0 30.1 37.5 39.7 38.2 37.3 39.7 45.4 41.3 37.7 37.3
NQ 49.3 64.8 62.3 62.8 68.9 73.5 68.7 70.4 73.7 70.2 75.3 67.3 69.1
HotpotQA 67.8 56.6 46.8 53.7 67.4 71.4 64.0 71.6 72.0 79.0 76.2 66.0 65.9
FiQA-2018 21.6 21.1 22.1 25.5 25.7 29.2 24.4 27.8 32.3 30.5 36.5 27.0 27.5
Touché-2020 50.4 45.6 40.9 48.7 42.9 43.7 32.6 44.4 50.1 50.2 50.6 45.5 45.0
CQADupStack 27.3 25.4 25.7 30.1 24.5 33.5 32.6 30.2 35.7 33.1 36.1 30.4 30.7
DBPedia 33.3 49.2 44.4 43.5 51.8 55.5 52.6 52.6 54.0 55.1 48.1 49.1 50.7
SCIDOCS 14.1 10.7 9.5 11.4 13.3 14.2 12.4 13.6 14.8 15.8 16.7 13.3 13.2
FEVER 71.4 82.5 78.1 82.4 85.8 88.1 78.5 89.2 89.5 89.9 79.5 83.2 84.3
Climate-FEVER 22.1 23.8 23.9 24.0 28.6 28.0 23.7 27.9 29.0 31.3 31.0 26.7 27.1
SciFact 62.0 41.0 43.0 42.9 56.4 59.6 53.0 57.1 59.9 61.2 68.0 54.9 54.2
@@
Supervised 47.2 69.4 68.5 66.7 69.9 68.5 70.7 70.8 73.7 71.3 69.1 67.8 69.8
Zero-shot 45.4 45.3 42.2 45.8 49.5 51.7 47.8 51.6 53.4 55.1 55.7 49.4 49.8
All 45.8 49.9 47.1 49.7 53.3 54.9 52.1 55.2 57.2 58.2 58.2 52.9 53.6